# Infinite-component $BF$ topological field theory: Nexus of fracton order, Toeplitz braiding, and non-Hermitian Amplification

### v2

@@ II Braiding in iBFBF theories
II.1 Multi-component $BF$ theories
In 3d Abelian topological orders, the topological excitations include both particles and loops; the $BF$ theory naturally captures the associated Aharonov–Bohm particle–loop braiding phase. The partition function of a multi-component $BF$ theory [^5] While a single-component $BF$ theory usually suffices for a given 3+13{+}1D spacetime, the multi-component generalization is essential for our construction of Toeplitz braiding. A related example is the charge–loop excitation symmetry in Ref. [123]. is
@@ II.1 Multi-component $BF$ theories
The integral $\int\delta\Omega\!\wedge\!\delta\gamma$ equals the linking number between $\Omega$ and $\gamma$. Thus, the Aharonov–Bohm phase between a particle and a loop is [123]
@@
This shows that particle–loop braiding phases are encoded in $K^{-1}$, directly analogous to particle–particle braiding in (2+1)D multi-component Chern–Simons theories.
@@ II.2 Construction of infinite-component $BF$ theories
As outlined in Sec. I, the iBFBF theory arises from coupling multi-component (3+1)D $BF$ theories of the form $S=\int\frac{A_{I,J}}{2\pi}\,b^{I}\wedge da^{J}$, where $b^{I}$ and $a^{J}$ denote compact $U(1)$ 2-form and 1-form gauge fields, respectively. Here $A$ is a square integer matrix, and repeated indices $I,J$ are summed over. By stacking $N$ layers of 3d topological orders described by these $BF$ theories and introducing interlayer $BF$ couplings that preserve translational symmetry along the stacking direction, we obtain a multi-component $BF$ theory characterized by an integer Toeplitz $K$ matrix that encodes the braiding statistics of topological excitations, as given in Eq. (2).
@@
Fig. 1 schematically illustrates this stacking and particle–loop braiding process: each layer is a 3-torus (blue cubes in Figs. 1(a)–(b)), the wavy lines represent interlayer $BF$ terms, and the stacking direction is the $w$ axis. Here, we introduce a useful quantity, $\Theta_{I,J}$, which encodes the spatial distribution of particle–loop braiding phases along the $w$ direction:
@@
where the two entries $1$’s are respectively located at the $I^{\text{th}}$ and $J^{\text{th}}$ components of $\mathbf{N}$ and $\mathbf{L}$ i.e., $N_{I^{\prime}}=\delta_{I,I^{\prime}},L_{J^{\prime}}=\delta_{J,J^{\prime}}$. Accordingly, the braiding between a loop at the $w_{1}$ boundary and a particle at the $w_{2}$ boundary [Fig. 1(a)] is encoded in the lower-left elements of $K^{-1}$ as well as the matrix $\Theta$ [Fig. 1(c)], while the reverse configuration [Fig. 1(b)] corresponds to the upper-right elements [Fig. 1(d)].
Before analyzing specific iBFBF theories, we emphasize the necessity of considering non-diagonal Toeplitz $K$ matrices in Eq. (2). Conventional multi-component $BF$ theories possess a “relabeling redundancy” [123]: if two different $K$ matrices denoted as $K$ and $K^{\prime}$ of the same size $N\times N$, are related by two independent $GL(N,\mathbb{Z})$ transformations $\Omega,W$ ($|\det\Omega|=|\det W|=1$, $W_{I,J},\Omega_{I,J}\in\mathbb{Z}$) as $K^{\prime}=\Omega KW^{\mathsf{T}}$, and the particle and loop charge vectors transform as $\mathbf{N}^{\prime}=W\mathbf{N}$ and $\mathbf{L}^{\prime}=\Omega\mathbf{L}$. Then we obtain identical braiding phases, $2\pi{\mathbf{N}^{\prime}}^{\mathsf{T}}{{K}^{\prime}}^{-1}\mathbf{L}^{\prime}=2\pi\mathbf{N}^{\mathsf{T}}{K}^{-1}\mathbf{L}$. Since any integer matrix can be brought to its Smith normal form via $GL(N,\mathbb{Z})$ transformations, one can normally decouple mutual $BF$ terms by such diagonalization and study the diagonalized $K$ matrix. However, by noting that $I,J$ indices now label the $w$ coordinates for iBFBF theories, preserving locality along the stacking direction restricts the allowed $GL(N,\mathbb{Z})$ transformations for Toeplitz $K$ matrices [Eq. (2)]: $\Omega$ and $W$ must be quasi-diagonal, with nonzero elements only clustered near the main diagonal.
@@
III iBFBF theory with Hatano-Nelson type $K$ matrix
@@
In this section and following section, we demonstrate these features using two representative examples: the iBFBF theory with a HN–type $K$ matrix and that with a nSSH–type $K$ matrix. Both of them arise from coupling multi-component $BF$ theories [Eq. (2)] with either $A=n$ or $A=\begin{pmatrix}A_{11}&A_{12}\\[2.0pt] A_{21}&A_{22}\end{pmatrix},\qquad\det A\neq 0.$ For a $2\times 2$ integer matrix $A$ of this form, its Smith normal form is $\begin{pmatrix}d_{1}&0\\[2.0pt] 0&d_{2}\end{pmatrix}$, where $d_{1}=\gcd(A_{11},A_{12},A_{21},A_{22})$ and $d_{2}=|A_{11}A_{22}-A_{12}A_{21}|/\gcd(A_{11},A_{12},A_{21},A_{22})$ [173]. Hence, the corresponding $BF$ theory effectively describes a $d_{1}\times d_{2}$ gauge theory.
@@ III.1 Toeplitz braiding and ZSMs–a concrete example
If $|b|>|n|$ together with $n\,\mathrm{rad}(n)\nmid b$ [^6] Here $\mathrm{rad}(n)$ denotes the product of the distinct prime factors of $n$. The condition $n\,\mathrm{rad}(n)\nmid b$ simply means that $b$ is not divisible by $n\,\mathrm{rad}(n)$, such that the braiding phase angle $\Theta_{I,J}$ defined in Eq. (9) is not trivially equal to $0\text{ mod }2\pi$., a nontrivial braiding statistical phase between the particle and the loop arises. From the equation for braiding phase (Eq. (8)) together with the expression for $K_{\mathrm{HN}}^{-1}$, only when a particle is placed on $w_{2}$ boundary and a loop is placed on $w_{1}$ boundary, the nontrivial braiding phase manifests. For example, the mutual braiding phase between a particle $\mathbf{N}=(0,\ldots,0,1)^{\mathsf{T}}$ and a loop $\mathbf{L}=(1,0,\ldots,0)^{\mathsf{T}}$ is $2\pi\frac{1}{n}\left(-\frac{b}{n}\right)^{N-1}$. In contrast, if the positions of the particle and the loop are exchanged, the nontrivial braiding statistical phase between them vanishes. For illustrative purpose, we depict the $\Theta_{I,J}$ matrix (defined in Eq. (9)) for $K_{\mathrm{HN}}$ specified by $n=2,b=3$ with system size $N=80$ in Fig. 2(a), clearly demonstrating this nonlocal braiding feature in the $w$ direction.
Figure 2: Braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{HN}}^{-1})_{I,J}$ (Panel (a)) and $\Theta_{I,J}=2\pi(K_{\mathrm{HN,PBC}}^{-1})_{I,J}$ (Panel (b)) encoded in $K_{\mathrm{HN}}$ and $K_{\mathrm{HN,PBC}}$ specified by $n=2,b=3$, where we select the system size $N=80$ for illustrative purpose. The lower-left nonzero elements in Panel (b) corresponds to the braiding phases $\Theta_{I,J}$ between test particles and loops with $w$ coordinates $I\sim N$, $J\sim 1$, which are close in the $w$ direction due to PBC.
To deepen our understanding of the origin of this nonlocal braiding statistics, we dig into the SVD of $K_{\mathrm{HN}}$ matrix. The SVD of $K_{\mathrm{HN}}$ is written as $K_{\mathrm{HN}}=\sum_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}=\sigma_{1}\mathbf{u}_{1}\mathbf{v}_{1}^{\mathsf{T}}+\sigma_{2}\mathbf{u}_{2}\mathbf{v}_{2}^{\mathsf{T}}+\cdots$, where the left singular modes $\{\mathbf{u}_{i}\}$ and the right singular modes $\{\mathbf{v}_{i}\}$ are mutually orthogonal, i.e., $\mathbf{u}_{i}^{\mathsf{T}}\mathbf{u}_{j}=\delta_{i,j},~\mathbf{v}_{i}^{\mathsf{T}}\mathbf{v}_{j}=\delta_{i,j}$, and $\{\sigma_{i}\}$ are the corresponding singular values.
If $|b|>|n|$, $K_{\mathrm{HN}}$ possesses the following left singular mode $\mathbf{u}_{1}$ located at the $w_{1}$ boundary and right singular mode $\mathbf{v}_{1}$ located at the $w_{2}$ boundary when $N$ is sufficiently large:
@@
The singular modes such as $\mathbf{u}_{1}$ and $\mathbf{v}_{1}$ that reside at the boundaries, decay exponentially into the bulk, and have singular values approaching zero in the thermodynamic limit $N\!\to\!\infty$ are referred to as ZSMs, as introduced in Sec. I.2. Left and right ZSMs are abbreviated LZSM and RZSM, respectively [^7] The “left” and “right” here refer to internal SVD sectors as is introduced in Sec. I.2, and are not related to the spatial distribution of the ZSMs (see Footnote 3). The LZSMs, RZSMs and the singular values discussed in this work may differ from the conventional definitions by a sign, which can be migrated among them without affecting the corresponding terms in SVD.. For large $N$, we can construct an approximate matrix $M_{\mathrm{HN}}$ from $\sigma_{1}$, $\mathbf{u}_{1}$ and $\mathbf{v}_{1}$ to model the large-size behavior of $K^{-1}$:
@@ III.2 General theory of Toeplitz braiding and ZSMs
Given a more general example, how can we understand the correspondence between ZSMs and nonlocal braiding statistics along the stacking direction in a more general framework? For more general iBFBF theories with Toeplitz $K$ matrices, the connection between nonlocal braiding statistics and ZSMs can also be understood from the SVD of the $K$ matrix. The SVD of an arbitrary N×N\ $\times N$ matrix $K$ (not necessarily Toeplitz) is written as $K=\sum_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\dagger}=\sigma_{1}\mathbf{u}_{1}\mathbf{v}_{1}^{\dagger}+\sigma_{2}\mathbf{u}_{2}\mathbf{v}_{2}^{\dagger}+\cdots$, where the left singular modes $\{\mathbf{u}_{i}\}$, right singular modes $\{\mathbf{v}_{i}\}$ and the corresponding singular values $\{\sigma_{i}\}$ of $K$ are determined by [171, 174] $K\mathbf{v}_{i}=\sigma_{i}\mathbf{u}_{i},~K^{\dagger}\mathbf{u}_{i}=\sigma_{i}\mathbf{v}_{i},\quad i=1,\ldots,\mathcal{N}$. All singular modes can be chosen to be mutually orthogonal, i.e., $\mathbf{u}_{i}^{\dagger}\mathbf{u}_{j}=\delta_{i,j},~\mathbf{v}_{i}^{\dagger}\mathbf{v}_{j}=\delta_{i,j}$. The singular values are nonnegative by definition, and for any real matrix the corresponding singular modes can be chosen to be real [^8] As the $K$ matrices in this work are real matrices, we use the transpose ($\mathsf{T}$) rather than the Hermitian conjugate ($\dagger$}) in all expressions related to the SVD of Toeplitz $K$ matrices hereafter.. Moreover, the LZSMs, RZSMs and the singular values in this paper may differ from the conventional definitions by a sign, which can be migrated among them without affecting the corresponding terms in SVD.
In the following, LZSMs are denoted by $\mathbf{u}_{1},~\mathbf{u}_{2},\cdots$ and the RZSMs are denoted by $\mathbf{v}_{1},~\mathbf{v}_{2},\cdots$. For finite system size $N$, we write $K=\sum_{i}\sigma_{i}\,\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}$, so its inverse is given by $K^{-1}=\sum_{i}\sigma_{i}^{-1}\,\mathbf{v}_{i}\mathbf{u}_{i}^{\mathsf{T}}$. If the $K$ matrix possesses ZSMs, and all the LZSMs and RZSMs, namely $\mathbf{u}_{j}$ and $\mathbf{v}_{j}$, are located at the opposite $w$ boundaries, the ZSMs and their exponentially small singular values dominate the contribution to upper-right or lower-left elements of $K^{-1}$.
More specifically, if all the LZSMs localize at $w_{1}$ boundary and all the RZSMs localize at $w_{2}$ boundary, then, the ZSMs and their exponentially small singular values dominate the contribution to lower-left elements of $K^{-1}$ (see Fig. 1(e) column 1); if all the LZSMs localize at $w_{2}$ boundary and all the RZSMs localize at $w_{1}$ boundary, then, the ZSMs and their exponentially small singular values dominate the contribution to upper-right elements of $K^{-1}$ (see Fig. 1(e) column 2). In both cases, the bulk modes contribute exponentially small values to lower-left or upper-right elements of $K^{-1}$. That is,
@@
In particular, certain $K$ matrices can support multiple LZSMs localized at both $w$ boundaries, with RZSMs exhibiting similar pattern (see Fig. 1(e) column 3). In this case, nonzero elements exist for both the upper-right and lower-left regions of $K^{-1}$. These elements, together with the corresponding particle-loop braiding phases, are also dominated by the contributions from ZSMs. Concrete examples of this case are presented in Sec. IV.
@@
It is noteworthy that the presence of ZSMs does not necessarily lead to nontrivial braiding statistics between particles and loops located at opposite $w$ boundaries, since ZSMs alone do not guarantee the existence of noninteger elements in the upper-right or lower-left elements of $K^{-1}$, which are required for nontrivial braiding statistics. If the upper-right and lower-left elements of $K^{-1}$ are all integers, the corresponding braiding phases $2\pi(K^{-1})_{I,J}$ trivially take values of multiples of $2\pi$, which indicates that the nontrivial particle-loop braiding phases vanish for particles and loops placed sufficiently far away in the $w$ direction. For example, if $n\,\mathrm{rad}(n)\mathrel{|}b$, the $(K_{\mathrm{HN}}^{-1})_{I,J}$ in Eq. (15) is integer for sufficiently large |IJ||I-J|. More concretely, $K_{\mathrm{HN}}$ with $n=2$, $b=4$ is an example where ZSMs are present but nonlocal braiding statistics along $w$ direction is absent. The braiding phase $\Theta_{I,J}=\pi\,\delta_{I,J}$ is nonzero only for $I=J$, indicating that nontrivial braiding statistics is nontrivial only for particles and loops residing in the same layer. Nevertheless, such special cases are very dilute in the whole parameter space; hence, the appearance of ZSMs indeed indicates nonlocal braiding statistics along the $w$ direction in most cases.
@@
The same mathematical structure emerges in non-Hermitian directional amplification studies, which stem from researches in driven-dissipative systems [169, 170, 171, 172]. 1D driven-dissipative cavity arrays can display directional amplification of coherent inputs, with gains detected at the other boundary scale exponentially with system size. When the input is switched to the other end, the amplification effect is suppressed. Moreover, such exotic amplification effect absent in Hermitian systems can also be naturally understood from the singular value decomposition of the non-Hermitian Hamiltonian $H$. In systems with non-Hermitian directional amplification, the SVD spectrum of $H$ under OBC exhibits ZSMs, whose corresponding LZSMs $\{\mathbf{u}_{j}\}$ and RZSMs $\{\mathbf{v}_{j}\}$, localize exponentially at opposite edges. In the linear-response regime, the susceptibility $\chi=H^{-1}=\sum_{j}\sigma_{j}^{-1}\mathbf{v}_{j}\mathbf{u}_{j}^{\dagger}$ is thus dominated by contributions from these ZSMs, leading to exponentially large, unidirectional transmission between boundaries [171]. This SVD-based framework provides a mathematically transparent diagnosis to directional amplification of physical signals in driven-dissipative systems.
@@
Figure 3: Braiding statistics encoded in iBFBF theories with Hatano-Nelson-type (HN-type) $K$ matrices. (a1)–-(e1) and (a2)–(e2) demonstrate the Toeplitz braiding encoded in HN-type iBFBF theories, with the $K_{\mathrm{HN}}^{\prime}$ matrix specified by the parameters $n=5$, $b=1$, $c=5$ and $n=5$, $b=5$, $c=1$, respectively. (a3)–(e3) demonstrates a trivial iBFBF theory with $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5$, $b=3$ and $c=1$. (a1)–(a3) are the plots of the left singular modes of these matrices, while (b1)–(b3) are the plots of the right singular modes of these matrices. (c1)–(c3) are the matrix plots of the braiding phases $\Theta_{I,J}=2\pi({K_{\mathrm{HN}}^{\prime}}^{-1})_{I,J}$. For illustrative purpose, we take the system size NN to be 40. (d1)–(d3) and (e1)–(e3) demonstrate how the braiding phases $\Theta_{1,N}$ and $\Theta_{N,1}$ vary as the system size $N$ increases.

### v3

@@ II Braiding in iBFBF theories
II.1 Multicomponent $BF$ theories
In 3D Abelian topological orders, the topological excitations include both particles and loops; the $BF$ theory naturally captures the associated Aharonov–Bohm particle–loop braiding phase. The partition function of a multicomponent $BF$ theory [^5] While a single-component $BF$ theory usually suffices for a given 3+13{+}1D spacetime, the multicomponent generalization is essential for our construction of Toeplitz braiding. A related example is the charge–loop excitation symmetry in Ref. [122]. is
@@ II.1 Multicomponent $BF$ theories
The integral $\int\delta\Omega\!\wedge\!\delta\gamma$ represents the intersection number between $\Omega$ and $\gamma$, which is the linking number between $\omega$ and $\gamma$ by definition. Thus, the Aharonov–Bohm phase between a particle and a loop is [122]
@@
This shows that particle–loop braiding phases are encoded in $K^{-1}$, directly analogous to particle–particle braiding in (2+1)D multicomponent Chern–Simons theories.
@@ II.2 Construction of infinite-component $BF$ theories
As outlined in Sec. I, the iBFBF theory arises from coupling multicomponent (3+1)D $BF$ theories of the form $S=\int\frac{A_{I,J}}{2\pi}\,b^{I}\wedge da^{J}$, where $b^{I}$ and $a^{J}$ denote compact $U(1)$ 2-form and 1-form gauge fields, respectively. Here $A$ is a square integer matrix, and repeated indices $I,J$ are summed over. By stacking $N$ layers of 3D topological orders described by these $BF$ theories and introducing interlayer $BF$ couplings that preserve translational symmetry along the stacking direction, we obtain a multicomponent $BF$ theory characterized by an integer Toeplitz $K$ matrix that encodes the braiding statistics of topological excitations, as given in Eq. (2).
@@
Fig. 1 schematically illustrates this stacking and particle–loop braiding process: each layer is a 3-torus [blue cubes in Figs. 1(a) and 1(b)], the wavy lines represent interlayer $BF$ terms, and the stacking direction is the $w$ axis. Here, we introduce a useful quantity, $\Theta_{I,J}$, which encodes the spatial distribution of particle–loop braiding phases along the $w$ direction:
@@
where the two entries $1$’s are respectively located at the $I$th and $J$th components of $\mathbf{N}$ and $\mathbf{L}$ i.e., $N_{I^{\prime}}=\delta_{I,I^{\prime}},L_{J^{\prime}}=\delta_{J,J^{\prime}}$. Accordingly, the braiding between a loop at the $w_{1}$ boundary and a particle at the $w_{2}$ boundary [Fig. 1(a)] is encoded in the lower-left elements of $K^{-1}$ as well as the matrix $\Theta$ [Fig. 1(c)], while the reverse configuration [Fig. 1(b)] corresponds to the upper-right elements [Fig. 1(d)].
Before analyzing specific iBFBF theories, we emphasize the necessity of considering non-diagonal Toeplitz $K$ matrices in Eq. (2). Conventional multicomponent $BF$ theories possess a “relabeling redundancy” [122]: if two different $K$ matrices denoted as $K$ and $K^{\prime}$ of the same size $N\times N$, are related by two independent $GL(N,\mathbb{Z})$ transformations $\Omega,W$ ($|\det\Omega|=|\det W|=1$, $W_{I,J},\Omega_{I,J}\in\mathbb{Z}$) as $K^{\prime}=\Omega KW^{\mathsf{T}}$, and the particle and loop charge vectors transform as $\mathbf{N}^{\prime}=W\mathbf{N}$ and $\mathbf{L}^{\prime}=\Omega\mathbf{L}$, then we obtain identical braiding phases, $2\pi{\mathbf{N}^{\prime}}^{\mathsf{T}}{{K}^{\prime}}^{-1}\mathbf{L}^{\prime}=2\pi\mathbf{N}^{\mathsf{T}}{K}^{-1}\mathbf{L}$. Since any integer matrix can be brought to its Smith normal form via $GL(N,\mathbb{Z})$ transformations, one can normally decouple mutual $BF$ terms by such diagonalization and study the diagonalized $K$ matrix. However, by noting that $I,J$ indices now label the $w$ coordinates for iBFBF theories, preserving locality along the stacking direction restricts the allowed $GL(N,\mathbb{Z})$ transformations for Toeplitz $K$ matrices [Eq. (2)]: $\Omega$ and $W$ must be quasi-diagonal, with nonzero elements only clustered near the main diagonal.
@@
III iBFBF theory with Hatano-Nelson-type $K$ matrix
@@
In this section and following section, we demonstrate these features using two representative examples: the iBFBF theory with a HN–type $K$ matrix and that with a nSSH–type $K$ matrix. Both of them arise from coupling multicomponent $BF$ theories [Eq. (2)] with either $A=n$ or $A=\begin{pmatrix}A_{11}&A_{12}\\[2.0pt] A_{21}&A_{22}\end{pmatrix},\qquad\det A\neq 0.$ For a $2\times 2$ integer matrix $A$ of this form, its Smith normal form is $\begin{pmatrix}d_{1}&0\\[2.0pt] 0&d_{2}\end{pmatrix}$, where $d_{1}=\gcd(A_{11},A_{12},A_{21},A_{22})$ and $d_{2}=|A_{11}A_{22}-A_{12}A_{21}|/\gcd(A_{11},A_{12},A_{21},A_{22})$ [172]. Hence, the corresponding $BF$ theory effectively describes a $\mathbb{Z}_{d_{1}}\times\mathbb{Z}_{d_{2}}$ gauge theory.
@@ III.1 Toeplitz braiding and ZSMs–A concrete example
If $|b|>|n|$ together with $n\,\mathrm{rad}(n)\nmid b$, [^6] Here $\mathrm{rad}(n)$ denotes the product of the distinct prime factors of $n$. The condition $n\,\mathrm{rad}(n)\nmid b$ simply means that $b$ is not divisible by $n\,\mathrm{rad}(n)$, such that the braiding phase angle $\Theta_{I,J}$ defined in Eq. (9) is not trivially equal to $0\text{ mod }2\pi$. then a nontrivial braiding statistical phase between the particle and the loop arises. From the equation for braiding phase [Eq. (8)] together with the expression for $K_{\mathrm{HN}}^{-1}$, only when a particle is placed on $w_{2}$ boundary and a loop is placed on $w_{1}$ boundary, the nontrivial braiding phase manifests. For example, the mutual braiding phase between a particle $\mathbf{N}=(0,\ldots,0,1)^{\mathsf{T}}$ and a loop $\mathbf{L}=(1,0,\ldots,0)^{\mathsf{T}}$ is $2\pi\frac{1}{n}\left(-\frac{b}{n}\right)^{N-1}$. In contrast, if the positions of the particle and the loop are exchanged, then the nontrivial braiding statistical phase between them vanishes. For illustrative purpose, we depict the $\Theta_{I,J}$ matrix [defined in Eq. (9)] for $K_{\mathrm{HN}}$ specified by $n=2,b=3$ with system size $N=80$ in Fig. 2(a), clearly demonstrating this nonlocal braiding feature in the $w$ direction.
Figure 2: Braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{HN}}^{-1})_{I,J}$ (a) and $\Theta_{I,J}=2\pi(K_{\mathrm{HN,PBC}}^{-1})_{I,J}$ (b) encoded in $K_{\mathrm{HN}}$ and $K_{\mathrm{HN,PBC}}$ specified by $n=2,b=3$, where we select the system size $N=80$ for illustrative purpose. The lower-left nonzero elements in Panel (b) corresponds to the braiding phases $\Theta_{I,J}$ between test particles and loops with $w$ coordinates $I\sim N$, $J\sim 1$, which are close in the $w$ direction due to PBC.
To deepen our understanding of the origin of this nonlocal braiding statistics, we dig into the SVD of $K_{\mathrm{HN}}$ matrix. The SVD of $K_{\mathrm{HN}}$ is written as $K_{\mathrm{HN}}=\sum_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}=\sigma_{1}\mathbf{u}_{1}\mathbf{v}_{1}^{\mathsf{T}}+\sigma_{2}\mathbf{u}_{2}\mathbf{v}_{2}^{\mathsf{T}}+\cdots$, where the left singular modes $\{\mathbf{u}_{i}\}$ and the right singular modes $\{\mathbf{v}_{i}\}$ are mutually orthonormal, i.e., $\mathbf{u}_{i}^{\mathsf{T}}\mathbf{u}_{j}=\delta_{i,j},~\mathbf{v}_{i}^{\mathsf{T}}\mathbf{v}_{j}=\delta_{i,j}$, and $\{\sigma_{i}\}$ are the corresponding singular values.
If $|b|>|n|$, then $K_{\mathrm{HN}}$ possesses the following left singular mode $\mathbf{u}_{1}$ located at the $w_{1}$ boundary and right singular mode $\mathbf{v}_{1}$ located at the $w_{2}$ boundary when $N$ is sufficiently large:
@@
The singular modes such as $\mathbf{u}_{1}$ and $\mathbf{v}_{1}$ that reside at the boundaries, decay exponentially into the bulk, and have singular values approaching zero in the thermodynamic limit $N\!\to\!\infty$ are referred to as ZSMs, as introduced in Sec. I.2. Left and right ZSMs are abbreviated LZSM and RZSM, respectively. [^7] The “left” and “right” here refer to internal SVD sectors as is introduced in Sec. I.2, and are not related to the spatial distribution of the ZSMs (see Footnote 3). The LZSMs, RZSMs and the singular values discussed in this work may differ from the conventional definitions by a sign, which can be migrated among them without affecting the corresponding terms in SVD. For large $N$, we can construct an approximate matrix $M_{\mathrm{HN}}$ from $\sigma_{1}$, $\mathbf{u}_{1}$ and $\mathbf{v}_{1}$ to model the large-size behavior of $K^{-1}$:
@@ III.2 General theory of Toeplitz braiding and ZSMs
Given a more general example, how can we understand the correspondence between ZSMs and nonlocal braiding statistics along the stacking direction in a more general framework? For more general iBFBF theories with Toeplitz $K$ matrices, the connection between nonlocal braiding statistics and ZSMs can also be understood from the SVD of the $K$ matrix. The SVD of an arbitrary N×N\ $\times N$ matrix $K$ (not necessarily Toeplitz) is written as $K=\sum_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\dagger}=\sigma_{1}\mathbf{u}_{1}\mathbf{v}_{1}^{\dagger}+\sigma_{2}\mathbf{u}_{2}\mathbf{v}_{2}^{\dagger}+\cdots$, where the left singular modes $\{\mathbf{u}_{i}\}$, right singular modes $\{\mathbf{v}_{i}\}$ and the corresponding singular values $\{\sigma_{i}\}$ of $K$ are determined by $K\mathbf{v}_{i}=\sigma_{i}\mathbf{u}_{i},~K^{\dagger}\mathbf{u}_{i}=\sigma_{i}\mathbf{v}_{i},\quad i=1,\ldots,N$ [170, 173]. All singular modes can be chosen to be mutually orthonormal, i.e., $\mathbf{u}_{i}^{\dagger}\mathbf{u}_{j}=\delta_{i,j},~\mathbf{v}_{i}^{\dagger}\mathbf{v}_{j}=\delta_{i,j}$. The singular values are nonnegative by definition, and for any real matrix the corresponding singular modes can be chosen to be real [^8] As the $K$ matrices in this work are real matrices, we use the transpose ($\mathsf{T}$) rather than the Hermitian conjugate ($\dagger$}) in all expressions related to the SVD of Toeplitz $K$ matrices hereafter.. Moreover, the LZSMs, RZSMs and the singular values in this paper may differ from the conventional definitions by a sign, which can be migrated among them without affecting the corresponding terms in SVD.
In the following, LZSMs are denoted by $\mathbf{u}_{1},~\mathbf{u}_{2},\cdots$ and the RZSMs are denoted by $\mathbf{v}_{1},~\mathbf{v}_{2},\cdots$. For finite system size $N$, we write $K=\sum_{i}\sigma_{i}\,\mathbf{u}_{i}\mathbf{v}_{i}^{\mathsf{T}}$, so its inverse is given by $K^{-1}=\sum_{i}\sigma_{i}^{-1}\,\mathbf{v}_{i}\mathbf{u}_{i}^{\mathsf{T}}$. If the $K$ matrix possesses ZSMs, and all the LZSMs and RZSMs, namely $\mathbf{u}_{j}$ and $\mathbf{v}_{j}$, are located at the opposite $w$ boundaries, then the ZSMs and their exponentially small singular values dominate the contribution to upper-right or lower-left elements of $K^{-1}$.
More specifically, if all the LZSMs localize at $w_{1}$ boundary and all the RZSMs localize at $w_{2}$ boundary, then, the ZSMs and their exponentially small singular values dominate the contribution to lower-left elements of $K^{-1}$ [see Fig. 1(e), column 1]; if all the LZSMs localize at $w_{2}$ boundary and all the RZSMs localize at $w_{1}$ boundary, then, the ZSMs and their exponentially small singular values dominate the contribution to upper-right elements of $K^{-1}$ [see Fig. 1(e), column 2]. In both cases, the bulk modes contribute exponentially small values to lower-left or upper-right elements of $K^{-1}$. That is,
@@
In particular, certain $K$ matrices can support multiple LZSMs localized at both $w$ boundaries, with RZSMs exhibiting similar pattern [see Fig. 1(e), column 3]. In this case, nonzero elements exist for both the upper-right and lower-left regions of $K^{-1}$. These elements, together with the corresponding particle-loop braiding phases, are also dominated by the contributions from ZSMs. Concrete examples of this case are presented in Sec. IV.
@@
It is noteworthy that the presence of ZSMs does not necessarily lead to nontrivial braiding statistics between particles and loops located at opposite $w$ boundaries, since ZSMs alone do not guarantee the existence of noninteger elements in the upper-right or lower-left elements of $K^{-1}$, which are required for nontrivial braiding statistics. If the upper-right and lower-left elements of $K^{-1}$ are all integers, the corresponding braiding phases $2\pi(K^{-1})_{I,J}$ trivially take values of multiples of $2\pi$, which indicates that the nontrivial particle-loop braiding phases vanish for particles and loops placed sufficiently far away in the $w$ direction. For example, if $n\,\mathrm{rad}(n)\mathrel{|}b$, then the $(K_{\mathrm{HN}}^{-1})_{I,J}$ in Eq. (15) is integer for sufficiently large |IJ||I-J|. More concretely, $K_{\mathrm{HN}}$ with $n=2$, $b=4$ is an example where ZSMs are present but nonlocal braiding statistics along $w$ direction is absent. The braiding phase $\Theta_{I,J}=\pi\,\delta_{I,J}$ is nonzero only for $I=J$, indicating that nontrivial braiding statistics is nontrivial only for particles and loops residing in the same layer. Nevertheless, such special cases are very dilute in the whole parameter space; hence, the appearance of ZSMs indeed indicates nonlocal braiding statistics along the $w$ direction in most cases.
@@
The same mathematical structure emerges in non-Hermitian directional amplification studies, which stem from researches in driven-dissipative systems [168, 169, 170, 171]. One-dimensional driven-dissipative cavity arrays can display directional amplification of coherent inputs, with gains detected at the other boundary scale exponentially with system size. When the input is switched to the other end, the amplification effect is suppressed. Moreover, such exotic amplification effect absent in Hermitian systems can also be naturally understood from the singular value decomposition of the non-Hermitian Hamiltonian $H$. In systems with non-Hermitian directional amplification, the SVD spectrum of $H$ under OBC exhibits ZSMs, whose corresponding LZSMs $\{\mathbf{u}_{j}\}$ and RZSMs $\{\mathbf{v}_{j}\}$, localize exponentially at opposite edges. In the linear-response regime, the susceptibility $\chi=H^{-1}=\sum_{j}\sigma_{j}^{-1}\mathbf{v}_{j}\mathbf{u}_{j}^{\dagger}$ is thus dominated by contributions from these ZSMs, leading to exponentially large, unidirectional transmission between boundaries [170]. This SVD-based framework provides a mathematically transparent diagnosis to directional amplification of physical signals in driven-dissipative systems.
@@
Figure 3: Braiding statistics encoded in iBFBF theories with Hatano-Nelson-type (HN-type) $K$ matrices. Panels (a1)–(e1) and (a2)–(e2) demonstrate the Toeplitz braiding encoded in HN-type iBFBF theories, with the $K_{\mathrm{HN}}^{\prime}$ matrix specified by the parameters $n=5$, $b=1$, $c=5$ and $n=5$, $b=5$, $c=1$, respectively. Panels (a3)–(e3) demonstrate a trivial iBFBF theory with $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5$, $b=3$ and $c=1$. Panels (a1)–(a3) are the plots of the left singular modes of these matrices, while panels (b1)–(b3) are the plots of the right singular modes of these matrices. Panels (c1)–(c3) are the matrix plots of the braiding phases $\Theta_{I,J}=2\pi({K_{\mathrm{HN}}^{\prime}}^{-1})_{I,J}$. For illustrative purpose, we take the system size NN to be 40. Panels (d1)–(d3) and (e1)–(e3) demonstrate how the braiding phases $\Theta_{1,N}$ and $\Theta_{N,1}$ vary as the system size $N$ increases.